# Netan Dogra

### v1

@@ 4. $2$-descent for Selmer schemes: local aspects away from $2$
For $v$ a prime of $K$ not lying above $p$, we define $H^{1}_{f,T}(G_{K_{v}},T\otimes\mathbb{F}_{p})$ to be the image of $H^{1}_{f}(G_{K_{v}},T)$ in $H^{1}(G_{K_{v}},T/pT)$. We will sometimes abbreviate this to $H^{1}_{f}(G_{K_{v}},T\otimes\mathbb{F}_{p})$ (when the lattice $T$ is unambiguous) but it should be emphasised that this subspace (unlike the subspace $H^{1}_{\mathrm{ur}}(G_{K_{v}},T\otimes\mathbb{F}_{p})$) depends on $T$ and is not an invariant of the Galois module $T\otimes\mathbb{F}_{p}$ in general.
4.1. Properties of $H^{1}_{f}(G_{K_{v}},T)$
@@
    (15) $H^{1}_{f}(G_{K_{v}},T\otimes\mathbb{F}_{p})\subset\mathrm{Ker}(H^{1}_{f}(G_{K_{v}},T\otimes\mathbb{F}_{p})\to H^{2}(G_{K_{v}},T)[p])$
@@
    $H^{1}_{f}(G_{K_{v}},T\otimes\mathbb{F}_{p})\subset H^{1}_{\mathrm{ur}}(G_{K_{v}},T_{p}).$
@@
Part 1 is immediate from $H^{1}_{f}(G_{K_{v}},T\otimes\mathbb{F}_{p})$ being defined to be a subspace of the image of $H^{1}(G_{K_{v}},T)$. For part 2, by definition $H^{1}_{f}(G_{K_{v}},T)$ consists of the classes whose image in $H^{1}(I_{v},T)^{G_{\mathbb{F}_{v}}}$ is torsion. Hence if the latter is torsion-free, then $H^{1}_{f}(G_{K_{v}},T)=H^{1}_{\mathrm{ur}}(G_{K_{v}},T)$. Since the image of $H^{1}_{\mathrm{ur}}(G_{K_{v}},T)$ in $H^{1}(G_{K_{v}},T\otimes\mathbb{F}_{p})$ is contained in $H^{1}_{\mathrm{ur}}(G_{K_{v}},T\otimes\mathbb{F}_{p})$, it is hence enough to show that $H^{1}(I_{v},T)^{G_{\mathbb{F}_{v}}}$ is torsion-free when $I_{v}$ acts unipotently. Since $I_{v}$ acts unipotently, we have
@@
where $I_{v}^{(p)}$ is the maximal pro-p quotient of $I_{v}$. Since this is isomorphic, as a group with an outer action of $G_{\mathbb{F}_{v}}$, to $\mathbb{Z}_{p}(1)$, we have an isomorphism of $G_{\mathbb{F}_{v}}$-modules
@@
We will write $H^{1}_{f}(G_{K_{v}},J[2])$ and $H^{1}_{f}(G_{K_{v}},\wedge^{2}J[2])$ to mean the image of $H^{1}_{f}(G_{K_{v}},T_{2}J)$ and $H^{1}_{f}(G_{K_{v}},\wedge^{2}T_{2}J)$ respectively.
@@
$H^{1}_{f}(G_{K},J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K}^{\times}\otimes\mathbb{F}_{2})$
@@
$H^{1}_{f}(G_{K},\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
@@
$H^{1}_{\mathrm{ur}}(K_{v},\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
@@ 5. $2$-descent for Selmer schemes: local aspects at $2$
To prove Theorem 1, it will be necessary to describe $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ as a subspace of $\mathbb{Q}_{2,f}^{(2),\times}\otimes\mathbb{F}_{2}$. Recall that by definition $H^{1}_{f}(G_{\mathbb{Q}_{2}},\wedge^{2}J[2])$ is the image of $H^{1}_{f}(G_{\mathbb{Q}_{2}},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$.We aim to achieve this using the nonabelian $(x-T)$ map. The idea behind our approach is as follows. We can calculate the dimension of the image of $H^{1}_{f}(G_{\mathbb{Q}_{2}},\wedge^{2}T_{2}J)$ in $H^{1}(G_{\mathbb{Q}_{2}},\wedge^{2}J[2])$, hence to describe the image it is enough to construct enough classes in $H^{1}(G_{\mathbb{Q}_{2}},\wedge^{2}J[2])$ which have crystalline lifts. We do this by using $\mathbb{Q}_{2}$ points of $X$ to construct crystalline classes in a larger Galois cohomology set, and try to give conditions for these to ‘induce’ crystalline classes with values in $\wedge^{2}J[2]$.
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\infty,b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(G_{\mathbb{Q}_{2}},\Pi_{2})$, i.e. an element of $H^{1}(G_{\mathbb{Q}_{2}},\Pi_{2})$ whose image in $H^{1}(G_{\mathbb{Q}_{2}},U_{2}(\mathbb{Q}_{2}))$ lies in the image of the subset $H^{1}_{f}(G_{\mathbb{Q}_{2}},U_{2}(\mathbb{Q}_{2}))$ in the sense of Kim [Kim09]. Furthermore, if $z-b$ lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(G_{\mathbb{Q}_{2}},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
Now suppose we have another class $\gamma\in H^{1}_{f}(G_{\mathbb{Q}_{2}},\Pi_{2})$ such that the image of $\gamma$ in $H^{1}(G_{\mathbb{Q}_{2}},T_{2}J)$ is equal to that of $j_{\Pi_{2},b}(z)$. Then we may twist $\Pi_{2}$ by a cocycle defining $\gamma$ to get a new group $\Pi_{2}^{(\gamma)}$ which is still a central extension of $T_{2}J$ by $\wedge^{2}T_{2}J$, and under the bijection between $H^{1}(G_{\mathbb{Q}_{2}},\Pi_{2})$ and $H^{1}(G_{\mathbb{Q}_{2}},\Pi_{2}^{(\gamma)})$ (see e.g. [Ser02, I.5.34]), $j_{\Gamma,b}(z)$ will be sent to a new element which we denote $\gamma^{-1}j_{\Gamma,b}(z)$. Then $\gamma^{-1}j_{\Gamma,b}(z)$ defines an element of $H^{1}_{f}(G_{\mathbb{Q}_{2}},\wedge^{2}T_{2}J)$. Now suppose furthermore that the image of $\gamma$ in $H^{1}(G_{\mathbb{Q}_{2}},\Gamma)$ is trivial. Then the image of $\gamma^{-1}j_{\Gamma,b}(z)$ in $H^{1}(G_{\mathbb{Q}_{2}},\wedge^{2}J[2])$ is equal to that of $j_{\Gamma,b}(z)$. Hence we have verified the existence of a crystalline lift of $j_{\Gamma,b}(z)$.
@@ 5.1. Integral mixed extensions
The issue with the idea outlined above is that the only way we have of constructing elements of $H^{1}_{f}(G_{\mathbb{Q}_{2}},\Pi_{2})$ is from $\mathbb{Z}_{2}$-points of the curve $X$. This means that the condition of finding a $\gamma$ with the same image in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ as $[z]-[b]$ is too restrictive. For this reason, we replace $\Pi_{2}$ with a larger group for which it is easier to construct elements.
@@
Now let $\Pi_{n}$ be the maximal pro-$p$ $n$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}}-\infty,b)$. Let $T_{i}$ denote the $i$th graded piece of the Iwasawa algebra $\mathbb{Z}_{p}[\![\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{X}}-\infty,b)]\!]$ with respect to the $I$-adic filtration, where $I$ is the augmentation ideal. Define
@@
$(c_{1},c_{2})\mapsto c_{1}\cup c_{2}.$
In particular, a pair $(c_{1},c_{2})\in H^{1}(G,T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2})$ lies in the image of $H^{1}(G,U(T_{0},T_{1},T_{2}))$ if and only if $c_{1}\cup c_{2}$ is zero in $H^{2}(G,T_{0}^{*}\otimes T_{2})$.
@@
Suppose that $f$ is irreducible over $K$. Let $i_{f}:D_{f}\hookrightarrow H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$ under the right and left multiplication maps, and let $i:D\hookrightarrow H^{1}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$. Let $E_{f}(T_{2}J)$ be a crystalline extension
@@
$\iota_{M_{1},*}:H^{1}(G_{\mathbb{Q}_{2}},M_{1})\to H^{1}(G_{\mathbb{Q}_{2}},E(T_{2}J)).$
If $M$ is moreover crystalline, we can play the same game instead using the map $\iota_{E,f}$ to obtain a class in $H^{1}_{f}(G_{\mathbb{Q}_{2}},E_{f}(T_{2}J))$. We have the following obtain the following characterisation of $H^{1}_{f}(G_{\mathbb{Q}_{2}},E_{f}(T_{2}J))$.
@@
If $c_{1}\in H^{1}_{f}(G_{\mathbb{Q}_{2}},T_{2}J)$ and $c_{2}\in D_{f}$ have the property that there are crystalline mixed extensions $M_{1},\ldots,M_{n}$ such that
@@
Since $c_{1}\cup c_{2}=\sum\pi_{1*}(c_{1})\cup\pi_{2*}(c_{2})=0$, there is an integral mixed extension $M$ of $c_{1}$ and $c_{2}$. Then $[M]-\sum[M_{i}]$ is an element of $H^{1}(G_{\mathbb{Q}_{2}},E(T_{2}J))$ which lies in the image of $H^{1}(G_{\mathbb{Q}_{2}},I^{2}/I^{3})$, by the exact sequence
$H^{1}(G_{\mathbb{Q}_{2}},I^{2}/I^{3})\to H^{1}(G_{\mathbb{Q}_{2}},E(T_{2}J))\to H^{1}(G_{\mathbb{Q}_{2}},T_{2}J)\otimes D.$
In particular, we may twist $M$ by a class $\xi$ in $H^{1}(G_{\mathbb{Q}_{2}},I^{2}/I^{3})$ to obtain a mixed extension $M^{\xi}$ with the property that $[M]=\sum[M_{i}]$ in $H^{1}(G_{\mathbb{Q}},E(T_{2}J))$. In particular, since all the $M_{i}$ are in the image of $H^{1}_{f}(G_{\mathbb{Q}_{2}},E_{f}(T_{2}J))$, so is $M^{\xi}$, and hence $M^{\xi}$ is crystalline by Lemma 22. ∎
We obtain the following criterion for liftability which is amenable to computations. First recall that for any $\mathbb{F}_{2}$-vector space $M$ we may view $M$ as a subspace of $M^{\otimes 2}$ via the homomorphism $m\mapsto m^{\otimes 2}$. Let $Y:=X-\infty$ and let
@@
$\Theta_{3}:\mathbb{F}_{2}[Y(\mathbb{Z}_{2})\times Y(\mathbb{Z}_{2})]\to\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$
@@
$(P,Q)\mapsto(P-\infty)(Q-\infty)+(Q-\infty)^{2}.$
@@
Suppose $\Theta_{1}$ and $\Theta_{2}$ are surjective. Then for every $b,z\in X(\mathbb{Q}_{2})$ such that $z-b\in 4J(\mathbb{Q}_{2})$, the lift of $j_{\Gamma,b}(z)$ to $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is in the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$.
@@
where $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$ is embedded diagonally. Indeed it is clear that the image is contained in this subgroup, and to check it surjects onto it is enough to check it mod $2$ (i.e. that it surjects onto $M\otimes\mathbb{F}_{2}$). Modulo $2$ the target is an mixed extension with graded pieces $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2},(J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2})^{\otimes 2}/J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}\oplus\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$, and the mod $2$ map is an extension of $\Theta_{1}$ by $\Theta_{2}\oplus\Theta_{3}$. Hence, given $P,Q$ such that $P-Q\in 4\cdot J(\mathbb{Q}_{2})$, there are $\lambda_{i}$ in $\mathbb{Z}_{2}$ and $P_{i},Q_{i}$ in $Y(\mathbb{Z}_{2})$ such that

### v2

@@ 4. $2$-descent for Selmer schemes: local aspects away from $2$
For $v$ a prime of $K$ not lying above $p$, we define $H^{1}_{f,T}(K_{v},T\otimes\mathbb{F}_{p})$ to be the image of $H^{1}_{f}(K_{v},T)$ in $H^{1}(K_{v},T/pT)$. We will sometimes abbreviate this to $H^{1}_{f}(K_{v},T\otimes\mathbb{F}_{p})$ (when the lattice $T$ is unambiguous) but it should be emphasised that this subspace (unlike the subspace $H^{1}_{\mathrm{ur}}(K_{v},T\otimes\mathbb{F}_{p})$) depends on $T$ and is not an invariant of the Galois module $T\otimes\mathbb{F}_{p}$ in general.
4.1. Properties of $H^{1}_{f}(K_{v},T)$
@@
    (15) $H^{1}_{f}(K_{v},T\otimes\mathbb{F}_{p})\subset\mathrm{Ker}(H^{1}(K_{v},T\otimes\mathbb{F}_{p})\to H^{2}(K_{v},T)[p])$
@@
    $H^{1}_{f}(K_{v},T\otimes\mathbb{F}_{p})\subset H^{1}_{\mathrm{ur}}(K_{v},T\otimes\mathbb{F}_{p}).$
@@
Part 1 is immediate from $H^{1}_{f}(K_{v},T\otimes\mathbb{F}_{p})$ being defined to be a subspace of the image of $H^{1}(K_{v},T)$. For part 2, by definition $H^{1}_{f}(K_{v},T)$ consists of the classes whose image in $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})}$ is torsion. Hence if the latter is torsion-free, then $H^{1}_{f}(K_{v},T)=H^{1}_{\mathrm{ur}}(K_{v},T)$. Since the image of $H^{1}_{\mathrm{ur}}(K_{v},T)$ in $H^{1}(K_{v},T\otimes\mathbb{F}_{p})$ is contained in $H^{1}_{\mathrm{ur}}(K_{v},T\otimes\mathbb{F}_{p})$, it is hence enough to show that $H^{1}(I_{v},T)^{\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})}$ is torsion-free when $I_{v}$ acts unipotently. Since $I_{v}$ acts unipotently, we have
@@
where $I_{v}^{(p)}$ is the maximal pro-p quotient of $I_{v}$. Since this is isomorphic, as a group with an outer action of $\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})$, to $\mathbb{Z}_{p}(1)$, we have an isomorphism of $\operatorname{Gal}(\overline{\mathbb{F}}_{v}|\mathbb{F}_{v})$-modules
@@
We will write $H^{1}_{f}(K_{v},J[2])$ and $H^{1}_{f}(K_{v},\wedge^{2}J[2])$ to mean the image of $H^{1}_{f}(K_{v},T_{2}J)$ and $H^{1}_{f}(K_{v},\wedge^{2}T_{2}J)$ respectively.
@@
$H^{1}_{f}(K,J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K}^{\times}\otimes\mathbb{F}_{2})$
@@
$H^{1}_{f}(K,\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
@@
$H^{1}_{\mathrm{ur}}(K,\wedge^{2}J[2])\simeq\mathrm{Ker}(\mathcal{O}_{K_{f}^{(2)}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}\mathcal{O}_{K_{f}}^{\times}\otimes\mathbb{F}_{2})$
@@ 5. $2$-descent for Selmer schemes: local aspects at $2$
To prove Theorem 1, it will be necessary to describe $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ as a subspace of $\mathbb{Q}_{2,f}^{(2),\times}\otimes\mathbb{F}_{2}$. Recall that by definition $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$. We aim to achieve this using the nonabelian $(x-T)$ map. The idea behind our approach is as follows. We can calculate the dimension of the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, hence to describe the image it is enough to construct enough classes in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ which have crystalline lifts. We do this by using $\mathbb{Q}_{2}$ points of $X$ to construct crystalline classes in a larger Galois cohomology set, and try to give conditions for these to ‘induce’ crystalline classes with values in $\wedge^{2}J[2]$.
More precisely, let $b$ and $z$ be $\mathbb{Z}_{2}$ points of $X$ not reducing to the point at infinity, and let $\Pi_{2}$ denote the maximal $2$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{\mathbb{Q}}_{2}}-\{\infty\},b)$. Then $j_{\Pi_{2},b}(z)$ is an element of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$, i.e. an element of $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ whose image in $H^{1}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ lies in the image of the subset $H^{1}_{f}(\mathbb{Q}_{2},U_{2}(\mathbb{Q}_{2}))$ in the sense of Kim [Kim09]. Furthermore, if $z-b$ lies in $4\cdot J(\mathbb{Q}_{2})$, then $j_{\Gamma,b}(z)$ lifts to an element of $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$, which we shall also denote $j_{\Gamma,b}(z)$ (although if $J[4](\mathbb{Q}_{2})\neq 0$ then this lift is not unique).
Now suppose we have another class $\gamma\in H^{1}_{f}(G_{\mathbb{Q}_{2}},\Pi_{2})$ such that the image of $\gamma$ in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ is equal to that of $j_{\Pi_{2},b}(z)$. Then we may twist $\Pi_{2}$ by a cocycle defining $\gamma$ to get a new group $\Pi_{2}^{(\gamma)}$ which is still a central extension of $T_{2}J$ by $\wedge^{2}T_{2}J$, and under the bijection between $H^{1}(\mathbb{Q}_{2},\Pi_{2})$ and $H^{1}(\mathbb{Q}_{2},\Pi_{2}^{(\gamma)})$ (see e.g. [Ser02, I.5.34]), $j_{\Gamma,b}(z)$ will be sent to a new element which we denote $\gamma^{-1}j_{\Gamma,b}(z)$. Then $\gamma^{-1}j_{\Gamma,b}(z)$ defines an element of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$. Now suppose furthermore that the image of $\gamma$ in $H^{1}(\mathbb{Q}_{2},\Gamma)$ is trivial. Then the image of $\gamma^{-1}j_{\Gamma,b}(z)$ in $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is equal to that of $j_{\Gamma,b}(z)$. Hence we have verified the existence of a crystalline lift of $j_{\Gamma,b}(z)$.
@@ 5.1. Integral mixed extensions
The issue with the idea outlined above is that the only way we have of constructing elements of $H^{1}_{f}(\mathbb{Q}_{2},\Pi_{2})$ is from $\mathbb{Z}_{2}$-points of the curve $X$. This means that the condition of finding a $\gamma$ with the same image in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ as $[z]-[b]$ is too restrictive. For this reason, we replace $\Pi_{2}$ with a larger group for which it is easier to construct elements.
@@
Now let $\Pi_{n}$ be the maximal pro-$p$ $n$-unipotent quotient of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)$. Let $T_{i}$ denote the $i$th graded piece of the Iwasawa algebra $\mathbb{Z}_{p}[\![\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{X}}-\{\infty\},b)]\!]$ with respect to the $I$-adic filtration, where $I$ is the augmentation ideal. Define
@@
$(c_{1},c_{2})\mapsto c_{1}\cup c_{2},$
In particular, a pair $(c_{1},c_{2})\in H^{1}(G,T_{0}^{*}\otimes T_{1}\oplus T_{1}^{*}\otimes T_{2})$ lies in the image of $H^{1}(G,U(T_{0},T_{1},T_{2}))$ if and only if $c_{1}\cup c_{2}$ is zero in $H^{2}(G,T_{0}^{*}\otimes T_{2})$ with respect to the pairing defined by the natural map
$T_{0}^{*}\otimes T_{1}\otimes T_{1}^{*}\otimes T_{2}\to T_{0}^{*}\otimes T_{2}.$
@@
Suppose that $f$ is irreducible over $\mathbb{Q}_{2}$. Let $i_{f}:D_{f}\hookrightarrow H^{1}_{f}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$ under the right and left multiplication maps, and let $i:D\hookrightarrow H^{1}(\mathbb{Q}_{2},\mathrm{Hom}(T_{2}J,I^{2}/I^{3}))$ denote the image of $H^{1}(\mathbb{Q}_{2},T_{2}J)^{\oplus 2}$. Let $E_{f}(T_{2}J)$ be a crystalline extension
@@
$\iota_{M_{1},*}:H^{1}(\mathbb{Q}_{2},M_{1})\to H^{1}(\mathbb{Q}_{2},E(T_{2}J)).$
If $M$ is moreover crystalline, we can play the same game instead using the map $\iota_{E,f}$ to obtain a class in $H^{1}_{f}(\mathbb{Q}_{2},E_{f}(T_{2}J))$. We deduce the following characterisation of $H^{1}_{f}(\mathbb{Q}_{2},E_{f}(T_{2}J))$.
@@
This implies the following criterion for the existence of crystalline mixed extensions lifting crystalline classes in $H^{1}(\mathbb{Q}_{2},T_{2}J)$ and $D\subset H^{1}(\mathbb{Q}_{2},T_{2}J^{*}\otimes I^{2}/I^{3})$.
@@
If $c_{1}\in H^{1}_{f}(\mathbb{Q}_{2},T_{2}J)$ and $c_{2}\in D_{f}$ have the property that there are crystalline mixed extensions $M_{1},\ldots,M_{n}$ such that
@@
Since $c_{1}\cup c_{2}=\sum\pi_{1*}(c_{1})\cup\pi_{2*}(c_{2})=0$, there is an integral mixed extension $M$ of $c_{1}$ and $c_{2}$. Then $[M]-\sum[M_{i}]$ is an element of $H^{1}(\mathbb{Q}_{2},E(T_{2}J))$ which lies in the image of $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$, by the exact sequence
$H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})\to H^{1}(\mathbb{Q}_{2},E(T_{2}J))\to H^{1}(\mathbb{Q}_{2},T_{2}J)\otimes D.$
In particular, we may twist $M$ by a class $\xi$ in $H^{1}(\mathbb{Q}_{2},I^{2}/I^{3})$ to obtain a mixed extension $M^{\xi}$ with the property that $[M]=\sum[M_{i}]$ in $H^{1}(\mathbb{Q},E(T_{2}J))$. In particular, since all the $M_{i}$ are in the image of $H^{1}_{f}(\mathbb{Q}_{2},E_{f}(T_{2}J))$, so is $M^{\xi}$, and hence $M^{\xi}$ is crystalline by Lemma 22. ∎
We obtain the following criterion for liftability which is amenable to computations. First recall that for any $\mathbb{F}_{2}$-vector space $M$ we may view $M$ as a subspace of $M^{\otimes 2}$ via the homomorphism $m\mapsto m\otimes m$. Let $Y:=X-\{\infty\}$ and let
@@
$\Theta_{3}:\mathrm{Ker}(\Theta_{1})\to\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$
@@
$(P,Q)\mapsto(P-\{\infty\})(Q-\{\infty\})+(Q-\{\infty\})^{2}.$
@@
Suppose $\Theta_{1}$ and $\Theta_{2}\oplus\Theta_{3}$ are surjective. Then for every $b,z\in X(\mathbb{Q}_{2})$ such that $z-b\in 4J(\mathbb{Q}_{2})$, the lift of $j_{\Gamma,b}(z)$ to $H^{1}(\mathbb{Q}_{2},\wedge^{2}J[2])$ is in the image of $H^{1}_{f}(\mathbb{Q}_{2},\wedge^{2}T_{2}J)$.
@@
where $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$ is embedded diagonally. Indeed it is clear that the image is contained in this subgroup, and to check it surjects onto it is enough to check it mod $2$ (i.e. that it surjects onto $M\otimes\mathbb{F}_{2}$). The vector space $M\otimes\mathbb{F}_{2}$ is an extension of $J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$ by $(J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2})^{\otimes 2}/J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}\oplus\mathrm{Sym}^{2}J(\mathbb{Q}_{2})\otimes\mathbb{F}_{2}$, and the map
$\mathbb{F}_{2}[Y(\mathbb{Z}_{2})\times Y(\mathbb{Z}_{2})]\to M\otimes\mathbb{F}_{2}$
is an extension of $\Theta_{1}$ by $\Theta_{2}\oplus\Theta_{3}$. Hence, given $P,Q$ such that $P-Q\in 4\cdot J(\mathbb{Q}_{2})$, there are $\lambda_{i}$ in $\mathbb{Z}_{2}$ and $P_{i},Q_{i}$ in $Y(\mathbb{Z}_{2})$ such that